OpenStudy (jkbo):

Please help The sum of three consecutive odd numbers is 153. Which is the LARGEST number?  a.47 b.51 c.49 d.53

OpenStudy (shamil98):

49 + 51 + 53 = 153, so the answer is d. 53.

OpenStudy (mathstudent55):

First you need to assign variables to the unknown numbers. You have three unknown numbers here. Let's let x = the smallest of the three numbers.

OpenStudy (mathstudent55):

The next larger number is how much larger than x?

OpenStudy (shamil98):

all you need to do is look at the question. and think about it. You don't need to use algebra to solve this problem..

OpenStudy (jkbo):

thank you both I have one more question 4 · (7 · 9) = (4 · 7) · 9 is an example of which property? Commutative of Multiplication  Identity of Multiplication  Associative of Multiplication  Distributive  @shamil98 and @mathstudent55

OpenStudy (mathstudent55):

The commutative property means change of order. The identity of multiplication is 1 since a * 1 = a. Associative property is change of grouping. Distributive is a(b + c) = ab + ac. What does this look like to you?

OpenStudy (jkbo):

its identity @mathstudent55

OpenStudy (jkbo):

@mathstudent55 ?

OpenStudy (mathstudent55):

Does your problem mention anything about multiplying by 1? If not, then it's not identity.

OpenStudy (jkbo):

@mathstudent55 its associative

OpenStudy (mathstudent55):

correct

OpenStudy (jkbo):

thanks